### v4

# A Structural Characterization of the Hit Image in the Motivic Steenrod Algebra

Abstract.
The motivic hit problem asks for a minimal set of generators of $H^{*,*}(BV_{n};\mathbb{F}_{2})$ as a module over the motivic Steenrod algebra. Masaki Kameko [3] constructed, in the distinguished family of degrees $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$, a top layer spanned by the monotone translates of a single monomial $z_{k}$, and he showed that the Bockstein image in this layer is contained in a subspace generated by pairwise sums of these translates. In this paper, we work on the raw degree–$d$ component $N_{n}^{d,*}$, before passing to the quotient by the hit subspace. We therefore construct a local projection
$\vartheta:N_{n}^{d,*}\longrightarrow V,$
from the raw degree–$d$ component to Kameko’s $M_{1}$–summand $V=\langle\{\sigma(z_{k})|\ \sigma\in\mathrm{Mono}(k)\}\rangle$, together with a parity functional $\varepsilon:V\to\mathbb{F}_{2}$, and prove that
$\vartheta\bigl(A_{+}^{\sharp}(N_{n})\cap N_{n}^{d,*}\bigr)=\ker(\varepsilon).$
Thus parity gives an exact description of the local top-layer image of hit elements. As a consequence, any element whose local top-layer component has odd parity is non-hit. In particular, every odd-parity linear combination of the translates $\sigma(z_{k})$ represents a non-zero class in the motivic hit quotient.
Subsequently, a direct binary calculation shows that for
$n=2^{r}+1,\qquad k=n-4,\qquad r\geq 5,$
one has $\beta(d)>n$ for
$d=(n-1)(2^{k+1}-2)+k.$
Combined with Kameko’s non-hit theorem, this yields a new infinite family of counterexamples to the motivic Peterson-type conjecture, distinct from Kameko’s k=n3k=n-3 family. Our local parity criterion sharpens this conclusion by showing that every odd-parity linear combination of the translates $\sigma(z_{k})$ is non-hit in those degrees.
Finally, we prove a base-change invariance statement: the local parity criterion on $V$ (and hence its consequences, including the above family) persists over any algebraically closed field of characteristic 0.
Key words and phrases:
Motivic cohomology, motivic Steenrod algebra, hit problem, Peterson conjecture, parity functional
2020 Mathematics Subject Classification:
Primary 14F42, 55S10; Secondary 55S05, 55T15
1. Introduction
The classical hit problem asks for a minimal set of generators of the polynomial algebra
$P_{n}=\mathbb{F}_{2}[x_{1},\dots,x_{n}]$
as a module over the mod $2$ Steenrod algebra $A$; equivalently, it asks for a basis of the quotient
$QP_{n}^{d}:=P_{n}^{d}/(A^{+}(P_{n})\cap P_{n}^{d}),$
where $A^{+}\subset A$ denotes the ideal of positive-degree elements and $P_{n}^{d}$ is the homogeneous degree–$d$ part of $P_{n}$. This problem goes back to work of Peterson and Wood; see for instance the survey in Walker–Wood [7].
A key numerical invariant is the function
$\beta(d)=\min\Bigl\{s\in\mathbb{N}\,\Big|\,d=(2^{i_{1}}-1)+\cdots+(2^{i_{s}}-1),\ i_{j}\in\mathbb{N}\Bigr\},$
which already appears in Wood’s proof of the Peterson conjecture. Wood proved the following fundamental theorem [9].
Theorem 1.1 (Wood).
If $\beta(d)>n$, then $QP_{n}^{d}=0$.
Theorem 1.1 says that $\beta(d)>n$ is a sufficient condition for triviality of the hit quotient. Moreover, the condition is also necessary in the sense that the monomial $x_{1}^{2^{i_{1}}-1}\cdots x_{n}^{2^{i_{n}}-1}$ never lies in $A^{+}(P_{n})$. The interplay between the hit problem and the cohomology of $A$ via Singer’s transfer [5] connects this story to the classical Adams spectral sequence and the stable homotopy groups of spheres; see, e.g. [1], [4], [7, 8].
In motivic homotopy theory, Voevodsky introduced a motivic Steenrod algebra $A^{*,*}$ and reduced power operations in motivic cohomology over a base field of characteristic different from $2$ [6]. Prompted by the work of Dugger–Isaksen on the motivic Adams spectral sequence [2], it is natural to seek motivic analogues of the hit problem and of Peterson’s conjecture. Kameko initiated this study in [3], where he worked over the base field $\mathbb{C}$ and defined, for each $n\geq 1$, the bigraded ring
$M_{n}(\mathbb{C}):=H^{*,*}(BV_{n}/\mathbb{C};\mathbb{F}_{2})$
together with its $A^{*,*}$–module structure. He then formulated the motivic hit problem for the quotients
$QM_{n}^{d,*}(\mathbb{C}):=M_{n}(\mathbb{C})_{d,*}/(A^{>0,*}\cdot M_{n}(\mathbb{C})\cap M_{n}(\mathbb{C})_{d,*}),$
where $A^{>0,*}$ is the subspace of $A^{*,*}$ consisting of elements of positive total degree (topological plus weight). In direct analogy with Theorem 1.1, he proposed the following motivic version of Peterson’s conjecture.
Definition 1.2 (Motivic Peterson-type conjecture [3, Conj. 1.2]).
Assume the base field is $\mathbb{C}$. The motivic Peterson conjecture asserts that
$\beta(d)>n\quad\Longrightarrow\quad QM_{n}^{d,*}(\mathbb{C})=0$
for all $n,d$.
Kameko showed that this conjecture fails in general. For each $1\leq k<n$ he defined an explicit monomial $z_{k}\in M_{n}(\mathbb{C})$ and proved that $z_{k}$ is not in the image of the motivic Steenrod algebra. Moreover, when k=n3k=n-3 and $\alpha(n-2)\geq 3$ (where $\alpha(t)$ denotes the number of $1$’s in the binary expansion of $t$), he showed that $\beta(d)>n$ for the degree
$d=(n-1)\bigl(2^{k+1}-2\bigr)+k,$
so that Conjecture 1.2 fails in those degrees. More precisely, we recall:
Theorem 1.3 (Kameko [3, Thm. 1.3, Props. 1.4–1.5]).
Assume the base field is $\mathbb{C}$. For each $n\geq 1$ and $1\leq k<n$ there is a monomial $z_{k}\in M_{n}(\mathbb{C})$ of degree $d=(n-1)(2^{k+1}-2)+k$ such that $z_{k}\notin A^{>0,*}\cdot M_{n}(\mathbb{C})$. If in addition $\alpha(n-2)\geq 3$ and k=n3k=n-3, then $\beta(d)>n$, so that $QM_{n}^{d,*}(\mathbb{C})\neq 0$ despite $\beta(d)>n$.
Kameko’s proof is quite delicate. He passes to the quotient $N_{n}(\mathbb{C})=M_{n}(\mathbb{C})/(\tau)$ and constructs a family of monomials in degree $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$ that give a basis for the top layer in the relevant filtration. He then shows that only certain linear combinations can appear in the image of the Bockstein $Q_{0}$, and that $z_{k}$ itself is not among them.
The present paper strengthens Kameko’s top-layer analysis by working on the raw degree–$d$ component $N_{n}^{d,*}$, before passing to the quotient by the full hit subspace. The key point is that, in degree $d=k+2d_{1}$, the part detected by Kameko’s set
$V=\bigl\langle\pi(\sigma(z_{k}))\ \big|\ \sigma\in\mathrm{Mono}(k)\bigr\rangle\subset\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}$
should be viewed as a local top layer, rather than as a subspace of the final quotient by hits. Here and throughout, $\mathrm{Mono}(k)$ denotes the set of monotone injections
$\sigma:\{1,\dots,k\}\hookrightarrow\{1,\dots,n\}.$
Equivalently, $\mathrm{Mono}(k)$ may be identified with the set of kk–subsets of $\{1,\dots,n\}$ via $\sigma\mapsto\mathrm{Im}(\sigma)$. Once this distinction is made, Kameko’s containment [3] for the $Q_{0}$–image becomes an exact parity statement for the local image of hit elements.
Our main contribution is therefore structural: we construct a local projection
$\vartheta:N_{n}^{d,*}\longrightarrow V$
and an explicit parity functional $\varepsilon:V\to\mathbb{F}_{2}$ such that
$\vartheta\bigl(A_{+}^{\sharp}(N_{n})\cap N_{n}^{d,*}\bigr)=\ker(\varepsilon).$
This gives a complete description of which vectors occur as the $M_{1}$–part of hit elements, and it yields an effective odd-parity obstruction.
  • (1)
    Local parity criterion. We prove that, in degree $d=k+2d_{1}$, the local top-layer image of hit elements is exactly the codimension-one hyperplane $\ker(\varepsilon)\subset V$. Equivalently, an element of $N_{n}^{d,*}$ whose local top-layer component has odd parity cannot be hit. In particular, every odd-parity linear combination of the translates $\sigma(z_{k})$ is non-hit.
  • (2)
    Arithmetic consequence and refinement. A direct binary calculation shows that if $n=2^{r}+1$ and k=n4k=n-4 with $r\geq 5$, then $\beta(d)>n$ for
    $d=(n-1)(2^{k+1}-2)+k.$
    Combined with Kameko’s non-hit theorem [3], this gives a new infinite family of motivic Peterson counterexamples. The local parity criterion refines this by proving that every odd-parity linear combination of the translates $\sigma(z_{k})$ is non-hit in those degrees.
  • (3)
    Base change. We show that the local parity criterion on $V$ is stable under extension of algebraically closed base fields of characteristic 0, and hence all of its consequences persist under base change.
Our main structural result is the following.
Theorem 1.4 (Local parity criterion in the top layer).
For integers $n\geq 2$ and $1\leq k<n$, let $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$, and let $V$ be the $\mathbb{F}_{2}$–span of the monotone translates $\sigma(z_{k})$ in Kameko’s $M_{1}$–summand. There exist a linear map
$\vartheta:N_{n}^{d,*}\longrightarrow V$
and a nonzero linear functional $\varepsilon:V\to\mathbb{F}_{2}$ such that:
  1. (i)
    Exact local top-layer hit image:
    $\vartheta\bigl(A_{+}^{\sharp}(N_{n})\cap N_{n}^{d,*}\bigr)=\ker(\varepsilon).$
  2. (ii)
    Odd-parity obstruction: if $u\in N_{n}^{d,*}$ satisfies $\varepsilon(\vartheta(u))=1$, then $u\notin A_{+}^{\sharp}(N_{n})$.
  3. (iii)
    In particular, any odd-parity linear combination of the monomials $\sigma(z_{k})$ is non-hit and therefore represents a non-zero class in $QN_{n}^{d,*}(\mathbb{C})$ and in $QM_{n}^{d,*}(\mathbb{C})$.
Remark 1.5 (Why the local formulation matters).
The map $\vartheta$ is defined on the raw degree–$d$ piece $N_{n}^{d,*}$, before quotienting by $A_{+}^{\sharp}(N_{n})$. This is essential: if one first passes to a quotient that already modds out hit elements, then every hit class becomes zero and the top-layer information is lost. The theorem above isolates the correct statement: parity detects exactly the local top-layer parts of hit elements.
On the arithmetic side, we show by a direct binary calculation that for $r\geq 5$, the choice $n=2^{r}+1$ and k=n4k=n-4 satisfies $\beta(d)>n$ for
$d=(n-1)(2^{k+1}-2)+k.$
Together with Kameko’s non-hit theorem, this yields a new infinite family of motivic Peterson counterexamples, distinct from Kameko’s original case k=n3k=n-3. The local parity criterion strengthens this conclusion by showing that every odd-parity linear combination of the translates $\sigma(z_{k})$ represents a non-zero class in the corresponding motivic hit quotient.
Combining Theorem 1.3, the local parity criterion of Theorem 1.4, the direct binary calculation for the case $n=2^{r}+1$, k=n4k=n-4, and the base-change discussion in Section 4, we obtain the following.
Theorem 1.6.
Let $K$ be an algebraically closed field of characteristic 0. For each integer $r\geq 5$, set $n=2^{r}+1$ and k=n4k=n-4, and let $d=(n-1)(2^{k+1}-2)+k$. Then:
  1. (1)
    $\beta(d)>n$.
  2. (2)
    The motivic hit quotient $QM_{n}^{d,*}(K)$ is non-zero.
  3. (3)
    More precisely, any odd-parity linear combination of the monomials $\sigma(z_{k})\in M_{n}(K)_{d,*}$, with $\sigma\in\mathrm{Mono}(k)$, represents a non-zero class in $QM_{n}^{d,*}(K)$.
Thus, combining the local parity description of the top layer with the direct verification of the inequality $\beta(d)>n$ in the case $n=2^{r}+1$, k=n4k=n-4, we obtain an infinite family of motivic Peterson-type counterexamples over any algebraically closed field of characteristic 0, distinct from Kameko’s k=n3k=n-3 family [3].
The paper is organized as follows. In Section 2 we review the classical hit problem, the motivic Steenrod algebra, and Kameko’s description of the motivic cohomology of $BV_{n}$ and the motivic hit problem. Section 3 introduces the local top-layer projection, defines the parity functional on Kameko’s $M_{1}$–summand, and proves Theorem 1.4. Finally, Section 4 discusses the extension to arbitrary algebraically closed base fields of characteristic 0, proves the base-change invariance of the local parity criterion, and combines it with the relevant arithmetic input to establish Theorem 1.6.
2. Preliminaries
In this section we collect definitions and known results needed later. We mostly follow the notation of Kameko [3].
2.1. The classical hit problem and the Peterson conjecture
Let $P_{n}=\mathbb{F}_{2}[x_{1},\dots,x_{n}]$ be the graded polynomial algebra with $|x_{i}|=1$. Let $A$ be the mod $2$ Steenrod algebra, and let $A^{+}\subset A$ be the ideal of elements of positive degree. Following the standard convention, we define
$A^{+}(P_{n})=\{a\cdot v\mid a\in A^{+},\ v\in P_{n}\}.$
The hit problem asks for a basis of the quotient
$QP_{n}^{d}=P_{n}^{d}/(A^{+}(P_{n})\cap P_{n}^{d})$
in each degree $d$. The numerical invariant $\beta(d)$ introduced in the introduction is closely related to the function $\alpha(t)$ counting $1$’s in the binary expansion of $t$.
Proposition 2.1 (Kameko [3, Prop. 2.1]).
For positive integers $d,n$ one has $\beta(d)>n$ if and only if $\alpha(d+n)>n$.
Wood proved that if $\beta(d)>n$ then $QP_{n}^{d}=0$; see Theorem 1.1. The condition is also necessary in the sense that monomials of the form $x_{1}^{2^{i_{1}}-1}\cdots x_{n}^{2^{i_{n}}-1}$ are never hit.
2.2. Motivic cohomology of $BV_{n}$ and the motivic Steenrod algebra
Let $k$ be a field of characteristic not equal to $2$. Motivic cohomology with $\mathbb{F}_{2}$–coefficients is a bigraded ring $H^{p,q}(-;\mathbb{F}_{2})$ defined on smooth schemes over $k$; we refer to Voevodsky [6] for the construction. The motivic Steenrod algebra $A^{*,*}$ is a bigraded Hopf algebra acting on motivic cohomology, generated by a Bockstein $Q_{0}$ of bidegree $(1,0)$ and reduced power operations $\mathcal{P}^{a}$ of bidegree $(2a,a)$ for $a\geq 1$, subject to Adem and instability relations. Over fields of characteristic zero, the algebra $A^{*,*}$ and its action on $H^{*,*}(-;\mathbb{F}_{2})$ are described in detail in [2, 6].
Let $V_{n}=(\mathbb{Z}/2)^{n}$ and write $BV_{n}$ for the corresponding classifying space. Over $\mathbb{C}$, Kameko shows that there is an isomorphism of bigraded rings
(2.1) $M_{n}(\mathbb{C}):=H^{*,*}(BV_{n}/\mathbb{C};\mathbb{F}_{2})\cong\mathbb{F}_{2}[\tau,x_{1},\dots,x_{n},y_{1},\dots,y_{n}]/(x_{1}^{2}+\tau y_{1},\dots,x_{n}^{2}+\tau y_{n}),$
where $|\tau|=(0,1)$, $|x_{i}|=(1,1)$, and $|y_{i}|=(2,1)$ [3, Sec. 1]. The mod $2$ motivic Steenrod algebra $A^{*,*}$ is generated by $Q_{0}$ and the $\mathcal{P}^{a}$ with actions [3, Sec. 1]:
$Q_{0}(\tau)$ $=0,$ $\mathcal{P}^{a}(\tau)$ $=0$ $(a\geq 1),$
$Q_{0}(x_{i})$ $=y_{i},$ $\mathcal{P}^{a}(x_{i})$ $=0$ $(a\geq 1),$
$Q_{0}(y_{i})$ $=0,$ $\mathcal{P}^{1}(y_{i})$ $=y_{i}^{2},\quad\mathcal{P}^{a}(y_{i})=0$ $(a\geq 2),$
together with the Cartan formulas.
It is convenient to pass to the quotient
$N_{n}(\mathbb{C}):=M_{n}(\mathbb{C})/(\tau)\cong\Lambda(x_{1},\dots,x_{n})\otimes\mathbb{F}_{2}[y_{1},\dots,y_{n}],$
and to the quotient $A^{\sharp}:=A^{*,*}/(\tau)$, which is again generated by $Q_{0}$ and the reduced powers $\mathcal{P}^{a}$. We denote by $A^{\sharp}_{+}$ the subspace of elements of positive total degree (topological plus weight). We write $A^{\sharp}_{+}(N_{n})=\{a\cdot u\mid a\in A^{\sharp}_{+},\ u\in N_{n}\}$.
2.3. The motivic hit problem and the monomials $z_{k}$
Following Kameko, we define the motivic hit quotients as follows.
Definition 2.2 (Motivic hit quotients [3, Sec. 1]).
For each topological degree $d$ let $M_{n}(\mathbb{C})_{d,*}=\bigoplus_{q}H^{d,q}(BV_{n}/\mathbb{C};\mathbb{F}_{2})$. Set
$QM_{n}^{d,*}(\mathbb{C}):=M_{n}(\mathbb{C})_{d,*}/(A^{>0,*}\cdot M_{n}(\mathbb{C})\cap M_{n}(\mathbb{C})_{d,*}),$
where $A^{>0,*}\subset A^{*,*}$ is the subspace of elements of positive total degree. The problem of describing a basis for the $QM_{n}^{d,*}(\mathbb{C})$ is the motivic hit problem.
Lemma 2.3 (Passage modulo $\tau$ preserves hit elements).
Let
$q_{\tau}:M_{n}(K)\longrightarrow N_{n}(K)=M_{n}(K)/(\tau)$
be the quotient map. If
$u\in A^{>0,*}\cdot M_{n}(K)\cap M_{n}(K)_{d,*},$
then
$q_{\tau}(u)\in A^{\sharp}_{+}(N_{n}(K))\cap N_{n}(K)_{d,*}.$
Consequently, $q_{\tau}$ induces a surjective linear map
$QM_{n}^{d,*}(K)\longrightarrow QN_{n}^{d,*}(K).$
Proof.
Write
$u=\sum_{i}a_{i}\cdot v_{i}\qquad(a_{i}\in A^{>0,*},\ v_{i}\in M_{n}(K)).$
Passing to the quotient modulo $\tau$, we obtain
$q_{\tau}(u)=\sum_{i}\bar{a}_{i}\cdot q_{\tau}(v_{i}),$
where $\bar{a}_{i}$ denotes the image of $a_{i}$ in
$A^{\sharp}=A^{*,*}/(\tau).$
If $\bar{a}_{i}=0$, the corresponding term vanishes. If $\bar{a}_{i}\neq 0$, then $\bar{a}_{i}$ still has positive total degree, hence $\bar{a}_{i}\in A^{\sharp}_{+}$. Therefore
$q_{\tau}(u)\in A^{\sharp}_{+}(N_{n}(K)).$
This proves the first assertion. The induced surjection on hit quotients follows immediately from the definition. ∎
The quotient $N_{n}(\mathbb{C})$ is an $A^{\sharp}$–module. By Lemma 2.3, the quotient map modulo $\tau$ induces a surjective map on hit quotients; in fact, one has an isomorphism
$QN_{n}^{d,*}(\mathbb{C}):=N_{n}(\mathbb{C})_{d,*}/\bigl(A^{\sharp}_{+}(N_{n}(\mathbb{C}))\cap N_{n}(\mathbb{C})_{d,*}\bigr)\cong QM_{n}^{d,*}(\mathbb{C})/\tau\,QM_{n}^{d,*}(\mathbb{C}),$
so we can work interchangeably with $QM_{n}^{d,*}(\mathbb{C})$ and $QN_{n}^{d,*}(\mathbb{C})$.
Kameko defines the function $\beta(d)$ exactly as in the classical case. As in (2.1), there are canonical monomials
$x_{1}\cdots x_{n}\,y_{1}^{2^{i_{1}-1}-1}\cdots y_{n}^{2^{i_{n}-1}-1}\in M_{n}(\mathbb{C})_{d,*}$
which are never in $A^{>0,*}\cdot M_{n}(\mathbb{C})$ when $\beta(d)\leq n$. This motivates the motivic Peterson-type conjecture (Conjecture 1.2).
For each pair of integers $n,k$ with $1\leq k<n$, Kameko defines a special monomial $z_{k}\in M_{n}(\mathbb{C})$ by
(2.2) $z_{k}=x_{1}\cdots x_{k}\left(\prod_{j=1}^{k}y_{j}^{2^{k}-2^{k-j}-1}\right)\left(\prod_{j=k+1}^{n}y_{j}^{2^{k}-1}\right),$
which has topological degree
(2.3) $d=(n-1)\bigl(2^{k+1}-2\bigr)+k.$
He proves:
Proposition 2.4 (Kameko [3, Prop. 1.4]).
For any integers $n\geq 1$ and $1\leq k<n$, the monomial $z_{k}$ defined in (2.2) is not in $A^{>0,*}\cdot M_{n}(\mathbb{C})$.
The key idea is to analyze the image of $Q_{0}$ in the quotient $N_{n}(\mathbb{C})/H_{n}$ for a certain subspace $H_{n}\subset N_{n}(\mathbb{C})$, and to describe the top layer in terms of two sets of monomials $M_{0}$ and $M_{1}$ which we review next.
2.4. Weight vectors and the subspaces $F_{n}$, $G_{n}$, and $H_{n}$
We briefly recall Kameko’s weight notation. Let $z=x_{1}^{\varepsilon_{1}}\cdots x_{n}^{\varepsilon_{n}}y_{1}^{e_{1}}\cdots y_{n}^{e_{n}}$ be a monomial in $N_{n}(\mathbb{C})$. Write
$\varepsilon_{i}+2e_{i}=\sum_{j\geq 0}\alpha_{ij}(z)2^{j},\qquad\alpha_{ij}(z)\in\{0,1\},$
and set
$\alpha_{i}(z)=\sum_{j\geq 0}\alpha_{ij}(z),\qquad\omega_{j}(z)=\sum_{i=1}^{n}\alpha_{ij}(z).$
Thus $\alpha_{i}(z)$ measures the number of non-zero bits in the binary expansion of $\varepsilon_{i}+2e_{i}$ and $\omega_{j}(z)$ measures how many of the $\varepsilon_{i}+2e_{i}$ have a $1$ in the $2^{j}$ place. We collect these into finite sequences
$\alpha(z)=(\alpha_{1}(z),\dots,\alpha_{n}(z)),\qquad\omega(z)=(\omega_{0}(z),\omega_{1}(z),\dots).$
Kameko defines a subspace $G_{n}\subset Y_{n}=\mathbb{F}_{2}[y_{1},\dots,y_{n}]$ generated by:
  • the image $P^{+}(Y_{n})$ of the positive-degree part of the reduced power subalgebra $P\subset A^{\sharp}$, and
  • all monomials $y\in Y_{n}$ such that $(\omega_{1}(y),\dots,\omega_{k}(y))<(n-1,\dots,n-1)$ in the lexicographic order.
He then uses an isomorphism $\psi:Y_{n}\to P_{n}$ sending $y_{i}$ to $x_{i}$ and intertwining the actions of $P$ and $A$ to identify $Y_{n}/G_{n}$ with the quotient $P_{n}/F_{n}$ from the classical hit problem in a suitable degree [3, Sec. 4].
On the $N_{n}$ side, Kameko defines a subspace $H_{n}\subset N_{n}$ spanned by $A^{\sharp}_{+}(N_{n})$ and those monomials $z$ for which
$(\omega_{0}(z),\omega_{1}(z),\dots,\omega_{k}(z))<(k,n-1,\dots,n-1)$
in the lexicographic order. He shows that the quotient $(N_{n}/H_{n})_{d,*}$ decomposes as a direct sum of pieces
$\Lambda_{n}^{a}\otimes(Y_{n}/G_{n})^{2b}/\pi\bigl(Q_{0}(\Lambda_{n}^{a+1}\otimes Y_{n}^{2(b-1)})\bigr)\qquad(a+2b=d),$
see [3, Sec. 4]. We will only need the degree $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$. In this case, Kameko constructs two sets of monomials $M_{0}$ and $M_{1}$ in $\Lambda_{n}^{k}\otimes Y_{n}^{2d_{1}}$ and shows that their images in $\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}$ form a basis. Moreover, he proves that
$M_{1}=\{\sigma(z_{k})\mid\sigma\in\mathrm{Mono}(k)\},$
and that the image of $Q_{0}$ in this summand is spanned, modulo $M_{0}$, by pairwise sums $\sigma_{1}(z_{k})+\sigma_{2}(z_{k})$; see [3, Props. 4.1, 5.1, 5.3].
We now turn to a more detailed analysis of this top summand.
3. A parity functional on Kameko’s local top layer
In this section we work in degree $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$. Our goal is to define a projection onto the $M_{1}$–summand before quotienting by hit elements and then to identify the image of hit elements under that projection.
3.1. The local top-layer map and the parity functional
Fix integers $n\geq 2$ and $1\leq k<n$. Let $d_{1}=(n-1)(2^{k}-1)$ and $d=k+2d_{1}$. Since
$N_{n}^{d,*}=\bigoplus_{a+2b=d}\Lambda_{n}^{a}\otimes Y_{n}^{2b},$
there is a canonical projection
$\operatorname{pr}_{k}:N_{n}^{d,*}\longrightarrow\Lambda_{n}^{k}\otimes Y_{n}^{2d_{1}}$
onto the summand with exactly $k$ exterior generators.
Let
$\pi:\Lambda_{n}^{k}\otimes Y_{n}^{2d_{1}}\longrightarrow\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}$
be the quotient map induced by $Y_{n}\to Y_{n}/G_{n}$. By Kameko’s description recalled above, there are subsets $M_{0},M_{1}\subset\Lambda_{n}^{k}\otimes Y_{n}^{2d_{1}}$ such that $\pi(M_{0})\cup\pi(M_{1})$ is a basis of $\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}$ and
(3.1) $M_{1}=\{\sigma(z_{k})\mid\sigma\in\mathrm{Mono}(k)\}.$
Set
$U_{0}:=\langle\pi(m)\mid m\in M_{0}\rangle,\qquad V:=\langle\pi(m)\mid m\in M_{1}\rangle.$
Then
$\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}=U_{0}\oplus V.$
Let
$p_{M_{1}}:\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}\longrightarrow V$
be the projection with kernel $U_{0}$.
Definition 3.1 (Local top-layer projection).
Define the linear map
$\vartheta:=p_{M_{1}}\circ\pi\circ\operatorname{pr}_{k}:N_{n}^{d,*}\longrightarrow V.$
We call $\vartheta(u)$ the local $M_{1}$–component (or local top-layer component) of $u\in N_{n}^{d,*}$.
The point is that $\vartheta$ is defined directly on $N_{n}^{d,*}$. It records the $M_{1}$–part of an element after discarding only the lower-weight terms encoded by $G_{n}$ and the complementary basis block $M_{0}$; it does not mod out by hit elements.
Definition 3.2 (Parity functional on the local top layer).
List $\mathrm{Mono}(k)$ as $\{\sigma_{1},\dots,\sigma_{N}\}$ and set $m_{i}=\sigma_{i}(z_{k})$, so $M_{1}=\{m_{1},\dots,m_{N}\}$ with $N=\binom{n}{k}$. Every element $v\in V$ can be written uniquely as
$v=\sum_{i=1}^{N}c_{i}\,\pi(m_{i}),\qquad c_{i}\in\mathbb{F}_{2}.$
We define
$\varepsilon:V\to\mathbb{F}_{2},\qquad\varepsilon(v)=\sum_{i=1}^{N}c_{i}\pmod{2}$.
Thus $\varepsilon(v)$ records whether the number of basis vectors $\pi(m_{i})$ appearing in $v$ is even or odd.
3.2. A linear algebra lemma
Lemma 3.3.
Let $W=\mathbb{F}_{2}^{N}$ with basis $e_{1},\dots,e_{N}$, and let $E\subset W$ be the subspace spanned by all pairwise sums $e_{i}+e_{j}$ with $i\neq j$. Then
$E=\{w=(w_{1},\dots,w_{N})\in W\mid w_{1}+\cdots+w_{N}=0\}.$
In other words, $E$ is exactly the hyperplane of vectors with even parity.
Proof.
For each $i\geq 2$ the vector $e_{1}+e_{i}$ lies in $E$. The N1N-1 vectors $e_{1}+e_{i}$ for $i=2,\dots,N$ are linearly independent and all have coordinate sum 0, so they span a subspace of dimension N1N-1 in the even-parity hyperplane. Thus $E$ has dimension at least N1N-1.
Conversely, let $w=(w_{1},\dots,w_{N})\in W$ satisfy $\sum_{i}w_{i}=0$. Then
$w=\sum_{i=1}^{N}w_{i}e_{i}=\sum_{i=2}^{N}w_{i}(e_{i}-e_{1})=\sum_{i=2}^{N}w_{i}(e_{1}+e_{i}),$
since over $\mathbb{F}_{2}$ we have $e_{1}=e_{1}$ and $w_{1}=\sum_{i=2}^{N}w_{i}$ by the parity assumption. Thus $w$ is a linear combination of the $e_{1}+e_{i}$, so $w\in E$.
We conclude that $E$ is an N1N-1–dimensional subspace of the even-parity hyperplane and coincides with it. ∎
3.3. The local $Q_{0}$–image in the $M_{1}$–summand
We now translate Kameko’s description of the image of $Q_{0}$ into a parity statement for the local map $\vartheta$.
Proposition 3.4 (Kameko [3, Props. 4.1, 5.1, 5.3]).
Suppose $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$ and $1\leq k<n$. Then:
  1. (1)
    The set $\pi(M_{0}\cup M_{1})$ is a basis of
    $\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}.$
    In particular,
    $\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}=U_{0}\oplus V.$
  2. (2)
    For each monomial $z\in N_{n}^{d-1,*}$ with $\omega_{0}(z)=k+1$, the element
    $\pi\bigl(Q_{0}(z)\bigr)\in\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}$
    is a linear combination of $\pi(z^{\prime})$ with $z^{\prime}\in M_{0}$ and elements of the form $\pi(\sigma_{1}(z_{k})+\sigma_{2}(z_{k}))$ with $\sigma_{1},\sigma_{2}\in\mathrm{Mono}(k)$.
  3. (3)
    There exist monotone injections $\sigma_{1},\sigma_{2}\in\mathrm{Mono}(k)$ with $\sigma_{1}\neq\sigma_{2}$ (explicitly, the ones with $\sigma_{1}(1)=1,\sigma_{1}(2)=3,\dots,\sigma_{1}(k)=k+1$ and $\sigma_{2}(1)=2,\dots,\sigma_{2}(k)=k+1$) and a monomial $z\in N_{n}^{d-1,*}$ with $(\omega_{0}(z),\omega_{1}(z))=(k+1,n-2)$ and $\alpha_{i}(z)=k$ for all $i$, such that
    $p_{M_{1}}\pi\bigl(Q_{0}(z)\bigr)=\pi\bigl(\sigma_{1}(z_{k})+\sigma_{2}(z_{k})\bigr).$
We can now identify the local $Q_{0}$–image in $V$.
Proposition 3.5 (Parity description of the local $Q_{0}$–image).
With notation as above, identify $V$ with $W=\mathbb{F}_{2}^{N}$ by sending the basis $\pi(\sigma_{i}(z_{k}))$ to $e_{i}$. Then
$\vartheta\bigl(Q_{0}(N_{n}^{d-1,*})\bigr)=\ker(\varepsilon)\subset V.$
Equivalently, the local $M_{1}$–part of the $Q_{0}$–image is exactly the even-parity hyperplane.
Proof.
If $z\in N_{n}^{d-1,*}$ has $\omega_{0}(z)\neq k+1$, then $Q_{0}(z)$ lies in a summand $\Lambda_{n}^{a}\otimes Y_{n}^{2b}$ with $a\neq k$, hence $\operatorname{pr}_{k}(Q_{0}(z))=0$ and therefore $\vartheta(Q_{0}(z))=0$. Thus only monomials with $\omega_{0}(z)=k+1$ contribute.
For such $z$, Proposition 3.4(2) shows that $\pi(Q_{0}(z))$ is a linear combination of $\pi(M_{0})$ and pairwise sums $\pi(\sigma_{1}(z_{k})+\sigma_{2}(z_{k}))$. After applying $p_{M_{1}}$, we conclude that $\vartheta(Q_{0}(z))$ is a linear combination of vectors of the form $e_{i}+e_{j}$. Hence
$\vartheta\bigl(Q_{0}(N_{n}^{d-1,*})\bigr)\subseteq E,$
where $E\subset W$ is the span of all $e_{i}+e_{j}$.
Conversely, Proposition 3.4(3) provides one explicit edge $e_{i_{0}}+e_{j_{0}}$ in the image. Let $S_{n}$ act on $N_{n}$ by permuting the variables $x_{1},\dots,x_{n},y_{1},\dots,y_{n}$. This action commutes with $Q_{0}$, preserves the subspace $G_{n}\subset Y_{n}$, and permutes the basis elements $\pi(\sigma(z_{k}))$ of $V$. Moreover, $S_{n}$ preserves $M_{0}$, since the defining condition for membership in $M_{0}$ is that $\alpha_{i}(z)<k$ for some $i$, and this condition is invariant under permutations of the indices. Hence $U_{0}=\langle\pi(m)\mid m\in M_{0}\rangle$ is $S_{n}$–stable, so the projection
$p_{M_{1}}:\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}\longrightarrow V$
with kernel $U_{0}$ is $S_{n}$–equivariant. Hence, for every permutation $\tau\in S_{n}$,
$\vartheta\bigl(Q_{0}(\tau\cdot z)\bigr)=p_{M_{1}}\pi\operatorname{pr}_{k}\bigl(Q_{0}(\tau\cdot z)\bigr)=\tau\cdot p_{M_{1}}\pi\operatorname{pr}_{k}\bigl(Q_{0}(z)\bigr)=\tau\cdot\vartheta\bigl(Q_{0}(z)\bigr).$
Therefore all edges of the Johnson graph $J(n,k)$ occur in the image, because the orbit of the explicit pair $(\sigma_{1},\sigma_{2})$ under $S_{n}$ consists exactly of pairs of kk–subsets differing in one element.
Since the Johnson graph $J(n,k)$ is connected for $1\leq k\leq n-1$, any vector $e_{i}+e_{j}$ is a sum of edge vectors along a path from $i$ to $j$. Thus every pairwise sum $e_{i}+e_{j}$ lies in $\vartheta(Q_{0}(N_{n}^{d-1,*}))$, so
$E\subseteq\vartheta\bigl(Q_{0}(N_{n}^{d-1,*})\bigr).$
Combining both inclusions gives
$\vartheta\bigl(Q_{0}(N_{n}^{d-1,*})\bigr)=E.$
Finally, by Lemma 3.3, $E=\ker(\varepsilon)$. ∎
3.4. From $Q_{0}$ to the full hit image in the local top layer
We now extend Proposition 3.5 from the Bockstein to the full positive-degree part of the motivic Steenrod algebra.
Lemma 3.6 (Reduced-power words vanish under $\vartheta$).
Let $\mathcal{W}$ be the set of finite words in the generators $Q_{0},\mathcal{P}^{1},\mathcal{P}^{2},\dots$ of $A^{\sharp}$. Then $A^{\sharp}_{+}$ is spanned by the images of elements of $\mathcal{W}$. If $w\in\mathcal{W}$ contains some reduced power $\mathcal{P}^{c}$ with $c\geq 1$, then
$\vartheta(w\cdot u)=0$
for every $u\in N_{n}$ such that $w\cdot u\in N_{n}^{d,*}$. Consequently,
$\vartheta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)=\vartheta\bigl(Q_{0}(N_{n}^{d-1,*})\bigr).$
Proof.
Since $A^{\sharp}_{+}$ is spanned by the images of words in $Q_{0},\mathcal{P}^{1},\mathcal{P}^{2},\dots$, it is enough by linearity to consider an operator represented by a single word $w$.
Assume that $w$ contains at least one reduced power. The reduced powers act trivially on the exterior factor $\Lambda_{n}$ and act on $Y_{n}$ through the reduced-power subalgebra $P\subset A^{\sharp}$. Because the word $w$ contains a reduced power, the induced operator on the $Y_{n}$–factor lies in $P^{+}$. Hence, for every $u\in N_{n}$, the $Y_{n}$–part of $w\cdot u$ belongs to
$P^{+}(Y_{n})\subset G_{n}.$
The Bockstein $Q_{0}$ acts only on the exterior factor and does not alter the $Y_{n}$–part. Therefore, after applying the projection
$\operatorname{pr}_{k}:N_{n}^{d,*}\longrightarrow\Lambda_{n}^{k}\otimes Y_{n}^{2d_{1}},$
the result lies in
$\Lambda_{n}^{k}\otimes G_{n}^{2d_{1}},$
which is annihilated by
$\pi:\Lambda_{n}^{k}\otimes Y_{n}^{2d_{1}}\longrightarrow\Lambda_{n}^{k}\otimes(Y_{n}/G_{n})^{2d_{1}}.$
Hence $\vartheta(w\cdot u)=0$.
Finally, since $Q_{0}^{2}=0$, the only nonzero word containing no reduced power is the single letter $Q_{0}$. The final equality follows. ∎
We can now state the main structural theorem of this section.
Theorem 3.7 (Theorem 1.4).
Let $n\geq 2$ and $1\leq k<n$, and let $d=k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$. Let $V$ be the local $M_{1}$–summand defined above, let $\vartheta:N_{n}^{d,*}\to V$ be the local top-layer projection from Definition 3.1, and let $\varepsilon:V\to\mathbb{F}_{2}$ be the parity functional from Definition 3.2. Then:
  1. (1)
    One has an equality of subspaces of $V$:
    $\vartheta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)=\ker(\varepsilon).$
  2. (2)
    If $u\in N_{n}^{d,*}$ satisfies $\varepsilon(\vartheta(u))=1$, then $u\notin A^{\sharp}_{+}(N_{n})$.
  3. (3)
    In particular, if
    $u=\sum_{\sigma\in S}\sigma(z_{k})\in N_{n}^{d,*}$
    for some non-empty subset $S\subset\mathrm{Mono}(k)$ and $|S|$ is odd, then $u\notin A^{\sharp}_{+}(N_{n})$.
Proof.
By Lemma 3.6,
$\vartheta\bigl(A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}\bigr)=\vartheta\bigl(Q_{0}(N_{n}^{d-1,*})\bigr).$
By Proposition 3.5, the latter space is exactly $\ker(\varepsilon)$. This proves (1).
For (2), if $u$ were hit, then $u\in A^{\sharp}_{+}(N_{n})\cap N_{n}^{d,*}$, and so part (1) would imply $\vartheta(u)\in\ker(\varepsilon)$, contradicting $\varepsilon(\vartheta(u))=1$.
For (3), note that if $u=\sum_{\sigma\in S}\sigma(z_{k})$, then $u\in\Lambda_{n}^{k}\otimes Y_{n}^{2d_{1}}$, its image under $\pi$ has no $M_{0}$–part, and therefore
$\vartheta(u)=\sum_{\sigma\in S}\pi(\sigma(z_{k})).$
Its parity is exactly $|S|$ modulo $2$. Hence odd cardinality implies $\varepsilon(\vartheta(u))=1$, and the claim follows from (2). ∎
Remark 3.8.
Theorem 3.7 identifies the image of hit elements after applying the local projection $\vartheta$. It does not assert that every even-parity linear combination of the monomials $\sigma(z_{k})$ is itself hit. The statement is weaker and sharper: every even-parity vector occurs as the local $M_{1}$–component of some hit element, while odd-parity vectors cannot occur in that way.
Corollary 3.9 (A large family of non-hit classes).
Let $S\subset\mathrm{Mono}(k)$ be a non-empty subset and set
$u_{S}:=\sum_{\sigma\in S}\sigma(z_{k})\in N_{n}^{d,*}.$
Let $\bar{u}_{S}$ be the image of $u_{S}$ in $QN_{n}^{d,*}(\mathbb{C})$ and hence in $QM_{n}^{d,*}(\mathbb{C})$. If $|S|$ is odd, then $\bar{u}_{S}$ is non-zero. In particular, each individual translate $\sigma(z_{k})$ represents a non-zero class, and any sum of an odd number of translates does as well.
Proof.
This is Theorem 3.7(3), interpreted in the hit quotient. ∎
4. Base change and motivic Peterson counterexamples over general fields
So far we have worked with the base field $\mathbb{C}$, as in Kameko’s original paper. We now explain why the constructions and counterexamples extend to any algebraically closed field $K$ of characteristic 0.
4.1. Motivic cohomology and the Steenrod algebra over general base fields
Let $K$ be an algebraically closed field of characteristic 0. Voevodsky constructed the mod $2$ motivic Steenrod algebra $A_{K}^{*,*}$ over any field of characteristic different from $2$ and described it in terms of generators and relations that depend only on the characteristic [6]. In particular, if $K$ and $\mathbb{C}$ are algebraically closed of characteristic 0, then $A_{K}^{*,*}\cong A_{\mathbb{C}}^{*,*}$ as bigraded Hopf algebras.
Similarly, for an elementary abelian 22–group $V_{n}=(\mathbb{Z}/2)^{n}$, the motivic cohomology ring $H^{*,*}(BV_{n}/K;\mathbb{F}_{2})$ is computed in terms of the cohomology of $B\mu_{2}$ and is independent of the choice of algebraically closed field of characteristic 0. Concretely, one has an isomorphism
(4.1) $M_{n}(K):=H^{*,*}(BV_{n}/K;\mathbb{F}_{2})\cong\mathbb{F}_{2}[\tau,x_{1},\dots,x_{n},y_{1},\dots,y_{n}]/(x_{1}^{2}+\tau y_{1},\dots,x_{n}^{2}+\tau y_{n}),$
with $|\tau|=(0,1)$, $|x_{i}|=(1,1)$ and $|y_{i}|=(2,1)$, and the action of $A_{K}^{*,*}$ is given by the same formulas as over $\mathbb{C}$; see, for example, [6, Sec. 9] together with [2]. In particular, there is a natural isomorphism of bigraded $A^{*,*}$–modules between $M_{n}(K)$ and $M_{n}(\mathbb{C})$, and likewise for their quotients $N_{n}(K)=M_{n}(K)/(\tau)$ and $N_{n}(\mathbb{C})$.
Remark 4.1.
No field embedding $K\hookrightarrow\mathbb{C}$ is needed here. For algebraically closed fields of characteristic 0, one has $\rho=[-1]=0\in H^{1,1}(\operatorname{Spec}K;\mathbb{F}_{2})$, just as for $K=\mathbb{C}$. In particular, the presentations (2.1) and (4.1) have the same defining relations, and the motivic Steenrod operations $Q_{0},\mathcal{P}^{a}$ act by the same generator formulas (together with Cartan) on the generators $\tau,x_{i},y_{i}$. Hence, using these explicit presentations, the assignment
$\tau\mapsto\tau,\qquad x_{i}\mapsto x_{i},\qquad y_{i}\mapsto y_{i}$
defines an isomorphism of bigraded rings $\Phi_{K}:M_{n}(K)\xrightarrow{\cong}M_{n}(\mathbb{C})$, and because the Steenrod action formulas on generators coincide, $\Phi_{K}$ is in fact an isomorphism of bigraded $A^{*,*}$–modules.
Using (4.1), we define the motivic hit quotients
$QM_{n}^{d,*}(K):=M_{n}(K)_{d,*}/(A^{>0,*}\cdot M_{n}(K)\cap M_{n}(K)_{d,*})$
exactly as in the $\mathbb{C}$–case, and similarly for the quotients $QN_{n}^{d,*}(K)$ of $N_{n}(K)$. The formulas and decompositions from Sections 2 and 3 make sense over $K$ and are preserved under the isomorphism $M_{n}(K)\cong M_{n}(\mathbb{C})$.
4.2. Extension of Kameko’s non-hit classes
We now show that Kameko’s non-hit monomials and our local parity criterion extend to arbitrary algebraically closed fields of characteristic 0.
Proposition 4.2.
Let $K$ be an algebraically closed field of characteristic 0. For any integers $n\geq 1$ and $1\leq k<n$, the monomial $z_{k}\in M_{n}(K)$ defined by (2.2) is not in $A^{>0,*}\cdot M_{n}(K)$.
Proof.
By Remark 4.1, there is an isomorphism of bigraded $A^{*,*}$–modules
$\Phi_{K}:M_{n}(K)\xrightarrow{\ \cong\ }M_{n}(\mathbb{C})$
sending $\tau,x_{i},y_{i}$ to the corresponding generators. In particular, $\Phi_{K}(z_{k})$ is the monomial $z_{k}$ in $M_{n}(\mathbb{C})$ considered by Kameko.
Suppose $z_{k}\in A^{>0,*}\cdot M_{n}(K)$. Applying $\Phi_{K}$, we obtain $\Phi_{K}(z_{k})\in A^{>0,*}\cdot M_{n}(\mathbb{C})$. But by Proposition 2.4, $\Phi_{K}(z_{k})$ is not in this image. This is a contradiction. Hence $z_{k}$ is not hit in $M_{n}(K)$. ∎
The same argument applies to odd-parity linear combinations of the translates $\sigma(z_{k})$.
Proposition 4.3.
Let $K$ be an algebraically closed field of characteristic 0, and let $n,k,d,M_{1},V,\varepsilon$ be as in Theorem 3.7. Then the conclusions of Theorem 3.7 and Corollary 3.9 hold over $K$: in particular, any odd-parity linear combination of the $\sigma(z_{k})$ represents a non-zero class in $QM_{n}^{d,*}(K)$.
Proof.
By Remark 4.1 there is an isomorphism of bigraded $A^{*,*}$–modules
$\Phi_{K}:M_{n}(K)\xrightarrow{\ \cong\ }M_{n}(\mathbb{C})$
sending $\tau,x_{i},y_{i}$ to the corresponding generators. Since $\Phi_{K}(\tau)=\tau$, it induces an isomorphism of $A^{\sharp}$–modules
$\Phi_{K}^{\sharp}:N_{n}(K)=M_{n}(K)/(\tau)\xrightarrow{\ \cong\ }M_{n}(\mathbb{C})/(\tau)=N_{n}(\mathbb{C}).$
Passing to hit quotients, we obtain an induced isomorphism
$\overline{\Phi}_{K}:QN_{n}^{d,*}(K)\xrightarrow{\ \cong\ }QN_{n}^{d,*}(\mathbb{C}).$
Under these identifications, the subspace $G_{n}$, the sets $M_{0},M_{1}$, the local top-layer summand $V$, the projection $\vartheta$, and the parity functional $\varepsilon$ are preserved, because they are defined by the same monomial formulas and the same weight conditions. Hence odd-parity classes over $K$ correspond exactly to odd-parity classes over $\mathbb{C}$.
Therefore, if $u=\sum_{\sigma\in S}\sigma(z_{k})$ has odd parity, then its image in $QN_{n}^{d,*}(\mathbb{C})$ is non-zero by Theorem 3.7. Since $\overline{\Phi}_{K}$ is an isomorphism, the class of $u$ is already non-zero in $QN_{n}^{d,*}(K)$. Finally, using $QN_{n}^{d,*}(K)\cong QM_{n}^{d,*}(K)/\tau\,QM_{n}^{d,*}(K)$, we conclude that $u$ represents a non-zero class in $QM_{n}^{d,*}(K)$ as well. ∎
4.3. Motivic Peterson counterexamples over $K$
We now combine Kameko’s arithmetic input in the case k=n3k=n-3 and a direct binary calculation in the case k=n4k=n-4 with the local parity statement over general base fields.
Theorem 4.4 (Motivic Peterson counterexamples for k=n3k=n-3 over $K$).
Let $K$ be an algebraically closed field of characteristic 0. Let $n\geq 9$ be an integer with $\alpha(n-2)\geq 3$, set k=n3k=n-3, and let $d=(n-1)(2^{k+1}-2)+k$. Then
  1. (1)
    $\beta(d)>n$.
  2. (2)
    The motivic hit quotient $QM_{n}^{d,*}(K)$ is non-zero.
In particular, Kameko’s motivic Peterson-type conjecture fails in these degrees over every such $K$.
Proof.
Assertion (1) is Kameko’s Proposition 1.5 [3], which states that if k=n3k=n-3 and $\alpha(n-2)\geq 3$, then $\beta(d)>n$.
For the non-vanishing of $QM_{n}^{d,*}(K)$, over $\mathbb{C}$ the monomial $z_{k}$ is non-hit by Theorem 1.3, hence $QM_{n}^{d,*}(\mathbb{C})\neq 0$. Since $M_{n}(K)\cong M_{n}(\mathbb{C})$ as $A^{*,*}$–modules, it follows that $QM_{n}^{d,*}(K)\neq 0$. Equivalently, one may invoke Proposition 4.2. ∎
Theorem 4.5 (New motivic Peterson counterexamples at distance $4$).
Let $K$ be an algebraically closed field of characteristic 0. Let $r\geq 5$ and set $n=2^{r}+1$ and k=n4k=n-4. Let $d=(n-1)(2^{k+1}-2)+k$ be the degree of $z_{k}$. Then:
  1. (1)
    $\beta(d)>n$.
  2. (2)
    The motivic hit quotient $QM_{n}^{d,*}(K)$ is non-zero.
  3. (3)
    More precisely, for any non-empty subset $S\subset\mathrm{Mono}(k)$ with $|S|$ odd, the element
    $u_{S}=\sum_{\sigma\in S}\sigma(z_{k})\in M_{n}(K)_{d,*}$
    represents a non-zero class in $QM_{n}^{d,*}(K)$.
Thus these degrees $(n,d)$ form an infinite family of counterexamples to the motivic Peterson-type conjecture over $K$. This family is distinct from the case k=n3k=n-3.
Proof.
We first prove $\beta(d)>n$. By Proposition 2.1, it is enough to show that $\alpha(d+n)>n$. Since k=n4k=n-4, we compute
$d+n$ $=(n-1)(2^{k+1}-2)+k+n$
$=(n-1)(2^{n-3}-2)+(n-4)+n$
$=(n-1)2^{\,n-3}-2(n-1)+2n-4$
$=(n-1)2^{\,n-3}-2.$
Now $n=2^{r}+1$, so $n-1=2^{r}$, and therefore
$d+n=2^{r}\cdot 2^{\,n-3}-2=2^{\,r+n-3}-2.$
Hence $d+n$ has binary expansion consisting of (r+n4)(r+n-4) ones followed by one zero, and thus
$\alpha(d+n)=r+n-4.$
Since $r\geq 5$, we have r+n4>nr+n-4>n, so $\alpha(d+n)>n$. Therefore $\beta(d)>n$ by Proposition 2.1.
For the non-vanishing and the explicit classes, note that
$d=k+2d_{1},\qquad d_{1}=(n-1)(2^{k}-1),$
so Theorem 3.7 applies in this degree. Hence any odd-parity sum
$u_{S}=\sum_{\sigma\in S}\sigma(z_{k})$
with $|S|$ odd represents a non-zero class in $QN_{n}^{d,*}(K)$ and therefore in $QM_{n}^{d,*}(K)$ by Proposition 4.3. Taking $S=\{\mathrm{id}\}$ gives, in particular, that $z_{k}$ itself is non-hit, so $QM_{n}^{d,*}(K)\neq 0$. ∎
Combining Theorems 4.4 and 4.5 proves Theorem 1.6 from the introduction.
Remark 4.6.
The family in Theorem 4.5 is genuinely distinct from Kameko’s original family [3]. In Kameko’s family one has k=n3k=n-3, while here k=n4k=n-4; the corresponding degrees $d$ differ, and the two families occupy different regions of the (n,d)(n,d)–plane.